Grade 12AIPMT

A BALL OF MASS 100g IS THROWN FROM THE GROUND WITH SPEED 20 m/s COMES BACK TO THE GROUND WITH SPEED 10 m/s . FIND THE WORK DONE BY THE AIR RESISTANCE.

RAGINI SEIWAL
8 Years agoGrade 12

1 Answer

Eshan
8 Years ago
Dear student,

Form law of conservation of energy, the work done by air resistance= loss in kinetic energy of the ball.

=\dfrac{1}{2}m(v_1^2-v_2^2)=\dfrac{1}{2}(0.1)(20^2-10^2)J=15J